OpenStudy (anonymous):

Solve 2 - 5x = x - 10 a) x = -3 b) x = -2 c) x = 2 d) x = 3

OpenStudy (anonymous):

x+5x=2+10 =>6x=12 =>x=2

OpenStudy (anonymous):

@zaibali.qasmi Solve -3(2x - 4) = 24 x = -2 x = 6 x = 2 x = -6

OpenStudy (anonymous):

multiply -3 in parenthesis.. -6x+12=24 =>6x=12-24 =>6x=-12 =>x=-2

OpenStudy (anonymous):

@zaibali.qasmi thank you so so much!